pytdestiny:

2v^2 − 7v + 4 = 0 Write your answers as integers, proper or improper fractions in simplest form, or decimals rounded to the nearest hundredth.

pytdestiny:

@snowflake0531

snowflake0531:

mmmm do you want to use the graphing method or use the quadratic equation?

pytdestiny:

quadratic equation

snowflake0531:

\[\frac{-b \pm \sqrt{b^2-4ac}}{2a} \\ \\ax^2+bx+c \\2x^2-7x+4\] \[ a=2 \\b=-7\\c=4\] Can you plug a, b, and c into the quadratic formula?

pytdestiny:

a= 2 b= -7 c= 4

snowflake0531:

yea... and plug that into \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

pytdestiny:

I'm confused

snowflake0531:

\[\frac{ -(-7) \pm \sqrt{-7)^2-4(2)(4)} }{ 2(2) }\]

pytdestiny:

14.43087m / p

pytdestiny:

@snowflake0531

snowflake0531:

uh wat can you just simplify that expression?

pytdestiny:

I put the equation in a calculator and that's what it gave me

snowflake0531:

Don't use a calculator.... use your brain o-o just simplify the numbers, multiply

pytdestiny:

7^2 = 49

snowflake0531:

\[\frac{ -(-7) \pm \sqrt{(-7)^2-4(2)(4)} }{ 2(2) } \\ \frac{ 7 \pm \sqrt{49-32} }{ 4 }\\\frac{ 7 \pm \sqrt{17} }{ 4 }\]

snowflake0531:

So then you have \[\frac{7+\sqrt17}{4}\\\frac{7-\sqrt17}{4}\]

pytdestiny:

8.03

snowflake0531:

sure find the decimal form of both

pytdestiny:

0.0803

snowflake0531:

i have no clue where you got your answers

1 attachment
snowflake0531:

1 attachment
snowflake0531:

So then you can round to 0.72 and 2.78